# Patherea: Cell Detection and Classification for the 2020s

### v1

@@ 8 Acknowledgment
The research was partially co-financed by funds from the Slovenian Research and Innovation Agency (ARIS), allocated to the research program groups P3-0054 (Pathology and Molecular Genetics) and P2-0214 (Computer Vision).
Appendix A Additional Results
A.1 Lizard
In Table 10 we report the results on the Lizard dataset (Graham et al., 2021) by using officially provided weights to initialize the ACFormer (Huang et al., 2023b) and PGT (Huang et al., 2023a) methods, without the need for re-training. No official training weights were provided for MCSpatNet (Abousamra et al., 2021) and FCRN (Xie et al., 2018).
We notice a slight drop in performance, when using the official weights. We report the best obtained Patherea and FCRN results after 100 training runs using using different random seeds. This gives an upper bound for the Lizard dataset. This is possible due to an efficient implementation of the Patherea approaches, which on average take around 2 hours to train on the Lizard dataset, in comparison with 2 days for MCSPatNet, ACFormer and PGT.
@@ A.1 Lizard
A.2 BRCA-M2C
We notice that MCSpatNet (Abousamra et al., 2021) performs much better with the official weights in comparison with 5x re-training (reported in Table 4) using the official code and training recipes. We report the best obtained Patherea and FCRN results after 100 training runs using different random seeds. We notice that we are able to reach the official performance of the MCSpatNet with one of the random seeds, which suggests that best performing MCSpatNet weights were selected for release. Similar improvement is noticeable for ACFormer (Huang et al., 2023b).
@@ A.2 BRCA-M2C
A.3 BCData
We report results on BCData (Huang et al., 2020) in Table 12. We directly report F1 scores from the original publication for competing methods SC-CNN (Sirinukunwattana et al., 2016), CSRNet (Li et al., 2018) and U-CSRNet (Huang et al., 2020). The official implementation code was not released. No details (or code) were provided on F1 score calculation. FCRN re-implementation (Xie et al., 2018) and Patherea results are reported using our F1 benchmarking code (Section 5).
@@ A.3 BCData
A.4 Qualitative Results
In Figures 5 and 6 we present qualitative results for Lizard (Graham et al., 2021) and BRCA-M2C (Abousamra et al., 2021) datasets. Lizard samples are collected from various public colorectal cancer datasets with various image size and quality. BRCA-M2C samples are collected and labeled from TCGA BRCA (Lingle et al., 2016).

### v2

@@ 8 Acknowledgment
The research was partially co-financed by funds from the Slovenian Research and Innovation Agency (ARIS), allocated to the research program groups P3-0054 (Pathology and Molecular Genetics) and P2-0214 (Computer Vision). The authors gratefully acknowledge the HPC RIVR consortium and EuroHPC JU for funding this research by providing computing resources of the HPC system Vega at the Institute of Information Science, Maribor, Slovenia.
Appendix A Patherea Dataset Annotation
The following cell types were annotated by pathologists in the Ki67-stained tissue sections:
1.
Positive tumor cells: tumor cells that stain positive for Ki67, indicating proliferating cells.
2.
Negative tumor cells: tumor cells that do not stain positive for Ki67, indicating resting (quiescent) cells.
3.
Other positive cells: Normal cells that stain positive for Ki67, such as endothelial cells, inflammatory cells, and mucosal epithelial cells, which are in proliferation phases.
4.
Other negative cells: Non-tumorous cells that do not stain positive for Ki67, for example, endothelial cells, inflammatory cells, non-neoplastic ducts and mucosal epithelial cells, which are not in proliferating phases.
Appendix B Additional Results
B.1 Lizard
In Table 10 we report the results on the Lizard dataset (Graham et al., 2021) by using officially provided weights to initialize the ACFormer (Huang et al., 2023b) and PGT (Huang et al., 2023a) methods, without the need for re-training. No official training weights were provided for MCSpatNet (Abousamra et al., 2021) and FCRN++.
We observe a slight performance drop when using official weights. We report the best Patherea-P2P and FCRN++ results from 100 training runs with different random seeds, providing an upper bound for the Lizard dataset. Our efficient Patherea-P2P implementation enables this extensive evaluation, requiring only 2 hours of training per-run on a single GPU, compared to 2 days on 4 GPUs for MCSPatNet, ACFormer, and PGT (B.7).
@@ B.1 Lizard
B.2 BRCA-M2C
We notice that MCSpatNet (Abousamra et al., 2021) performs much better with the official weights in comparison with 5x re-training (reported in Table 4) using the official code and training recipes. We report the best obtained Patherea-P2P and FCRN++ results after 100 training runs using different random seeds. We notice that we are able to reach the official performance of the MCSpatNet with one of the random seeds, which suggests that the best-performing MCSpatNet weights were selected for release. Similar improvement is noticeable for ACFormer (Huang et al., 2023b).
@@ B.2 BRCA-M2C
B.3 BCData
We report results on BCData (Huang et al., 2020) in Table 12. We directly report F1 scores from the original publication for competing methods SC-CNN (Sirinukunwattana et al., 2016), CSRNet (Li et al., 2018) and U-CSRNet (Huang et al., 2020). The official implementation code was not released. No details (or code) were provided on F1 score calculation. FCRN re-implementation (Xie et al., 2018) and Patherea-P2P results are reported using our F1 benchmarking code (Section 5).
@@ B.3 BCData
FCRN++ implementation is very similar to U-CSRNet (Huang et al., 2020), but achieves a lower performance. No implementation code or benchmarking details are available for U-CSRNet to verify the results.
B.4 Hungarian Matching - Evaluation
In Table 13 we report Patherea-LNET evaluation results using the Hungarian algorithm on the distance matrix, as opposed to the Boolean thresholded matrix (Section 5). Note that we only repeated the evaluation step on the same model predictions, as reported in Table 5.
| MCSpatNet | 0.753 | 0.706 | 0.189 | 0.430 | 0.520 | -3.8 |
| ACFormer | 0.794 | 0.754 | - | - | 0.387 | -3.9 |
| PGT | 0.766 | 0.735 | - | - | 0.375 | -3.2 |
| FCRN++ | 0.779 | 0.745 | 0.203 | 0.451 | 0.545 | -3.1 |
| ours (ViT) | 0.800 | 0.772 | 0.259 | 0.559 | 0.598 | -3.2 |
| ours (CNN) | 0.800 | 0.767 | 0.280 | 0.545 | 0.598 | -4.3 |
The results support the claims from Section 5 that the global objective in the Hungarian algorithm underestimates F1 scores due to some portion of wrongly matched predictions. The relative performance/ranking of the different approaches has not changed, as the error is roughly equally distributed (3-4%).
B.5 UNI Foundation Model
In Table 14 we report the results of using UNI pretrained foundation model (Chen et al., 2024). We used the same evaluation setup as for the results presented in Table 9. ViT-L backbone (Dosovitskiy et al., 2021) was used for Scratch, ImageNet, and UNI, in comparison to the ViT-B backbone used with Hibou experiments in Table 9 - as UNI pre-trained weights are not available for ViT-B. UNI family of models were not trained on any of the standard open datasets that are often used in labeled datasets, such as Lizard (Graham et al., 2021) and BRCA-M2C (Abousamra et al., 2021). In comparison to Hibou (Nechaev et al., 2024), only H&E slides were included in the UNI training dataset.
| Lizard | BRCA-M2C | LNET | |
| Scratch | 0.427 | 0.507 | 0.559 |
| ImageNet | 0.562 | 0.598 | 0.612 |
| UNI | 0.624 | 0.661 | 0.610 |
The results demonstrate a similar behaviour to Hibou results presented in Table 9. The results demonstrate that the boost on the Lizard and BRCA-M2C datasets is not from the TCGA data leakage, but rather an actual boost from the pre-trained model that our proposed method can effectively utilize.
B.6 Segmentation Model Comparison
In Table 15 we compare our Hibou foundation model approach from Table 3 against the CellViT++(Hörst et al., 2025) instance segmentation approach.
| CellViT++(SAM-H) | 0.622 | 0.327 | 0.787 | 0.678 | 0.331 | 0.422 | 0.528 |
| ours (ViT)† | 0.680 | 0.499 | 0.788 | 0.742 | 0.444 | 0.524 | 0.613 |
Our approach performed significantly better, despite being only trained on point-based annotations. In both cases, the backbone was frozen. For CellViT++, only a lightweight cell classification module was re-trained, while only ViT-Adapter (Chen et al., 2023) layers and prediction heads were trained for Patherea-P2P. Pre-trained segmentation backbone of CellViT++ improves the performance on rare classes (i.e., eosinophil, neutrophil) against some of the competing approaches in Table 3.
B.7 Runtime Analysis
We report an approximate training time per-fold on Patherea-LNET dataset in Table 16. The training time of our proposed Patherea-P2P method is multi-fold faster, regardless of the chosen backbone. Note that we didn’t include the pre-processing time needed to create intermediate representations for MCSpatNet (Abousamra et al., 2021) and FCRN++. We also only report the prompt fine-tuning training time for the PGT (Huang et al., 2023a) method.
The average inference time per image sample () is also reported. We used one of the Patherea-LNET test folds to compute the average inference time with a batch size set to 1 for all the approaches using a single Nvidia A100 40GB GPU. The proposed ViT-based approach is significantly faster in comparison with DETR-based approaches (ACFormer (Huang et al., 2023b), PGT (Huang et al., 2023a)), while the CNN backbone offers similar performance to density-based approaches (MCSpatNet (Abousamra et al., 2021), FCRN++.
| [ms] | ||
| MCSpatNet | 1d18h | 5 |
| ACFormer | 7d | 170 |
| PGT | 7d4h | 45 |
| FCRN++ | 4h | 9 |
| ours (ViT) | 2.1h | 21 |
| ours (CNN) | 1.5h | 13 |
B.8 Qualitative Results
In Figures 5, 6 and 7 we present qualitative results of different approaches and datasets. Figure 5 extends Patherea-LNET qualitative results from Figure 3, by adding the qualitative results from the competing approaches. Figures 6 and 7 show qualitative results of our proposed approach on Lizard (Graham et al., 2021) and BRCA-M2C (Abousamra et al., 2021). Lizard samples are collected from various public colorectal cancer datasets with various image sizes and resolutions. BRCA-M2C samples are collected and labeled from TCGA-BRCA (Lingle et al., 2016).